# The coisotropic embedding theorem for pre-symplectic vector spaces: an alternative proof

### v2

@@
The coisotropic embedding theorem for pre-symplectic vector spaces: an alternative proof
L. Schiavone1,2 [Uncaptioned image]
1Dipartimento di Matematica e Applicazioni Renato Caccioppoli, Università degli Studi di Napoli Federico II
2
e-mail: luca.schiavone@unina.it
@@ Acknowledgements
I acknowledge financial support from Next Generation EU through the project 2022XZSAFN – PRIN2022 CUP: E53D23005970006.
@@ Introduction
The Coisotropic Embedding Theorem (CET), originally proven by Mark Gotay in his seminal work on pre-symplectic manifolds [Got82], establishes that any pre-symplectic manifold can be embedded into a symplectic manifold in such a way that the original pre-symplectic manifold becomes a coisotropic submanifold. It has significant applications both in theoretical mathematics and applied fields, where it facilitates the study of constrained mechanical systems and reduction theory from various points of view, as we are going to recall in this introduction.
@@
In mathematical physics, Hamiltonian systems are geometrically formulated by means of symplectic manifolds (usually the cotangent bundle of a differential manifold, sometimes referred to as Phase Space) endowed with a Hamiltonian function [Arn89, AMRC78]. On the other hand, constrained Hamiltonian systems are usually described on submanifolds of the Phase Space (selected by the constraints) on which the symplectic form becomes degenerate (see [GNH77, CDI+24, CDI+22a] and references therein). Among various interesting physical systems, all gauge theories lie in this class of dynamical systems. The CET ensures that these constrained systems can still be analyzed within the framework of symplectic geometry by embedding the presymplectic manifold into a larger symplectic manifold [Got82, AMRC78, GS90].
@@
Recent advances in mathematical physics have highlighted the utility of the CET in various contexts. For instance, in [CDI+22a], the theorem is used to establish a Noether-like one-to-one correspondence between symmetries and constants of motion within pre-symplectic Hamiltonian systems. This correspondence is fundamental in understanding the conservation laws and invariant properties of constrained mechanical systems, thereby extending the classical Noether’s theorem to a broader context. Furthermore, we employed the CET to provide a Poisson structure on the space of solutions of certain field theories. For example, in [CDI+24, CDI+22b, CDI+22c], the theorem is used to pass from a pre-symplectic framework to a symplectic one, enabling the application of symplectic geometry and Poisson structures to the field theories considered, non-Abelian gauge theories and Palatini’s formulation of General Relativity in particular. Moreover, the theorem has been instrumental in addressing the inverse problem of the calculus of variations. In recent works [Sch24b, Sch24a], we used the CET to provide a solution for the inverse problem for a class of implicit differential equations. These examples illustrate how the theorem can be leveraged to construct variational principles for systems that are not explicitly variational. Additionally, it is worth mentioning that the CET has significant applications within the Geometric Quantization program [GS81]. Geometric Quantization is a procedure for constructing a quantum theory from a classical mechanical system, and requires a symplectic structure for the classical system as starting point. By embedding a pre-symplectic manifold into a symplectic one, the theorem facilitates the Quantization of constrained systems by ensuring that the necessary symplectic structure is available.
@@
The theorem also received several reformulations over the years. We mention the approach of V. Guilleming and S. Sternberg that proved an equivariant (with respect to the action of a Lie group) version of the theorem [GS90] and the approach of Y. G. Oh and J. S. Park [OP05] applied to the deformation of coisotropic submanifolds that makes even more transparent how the proof of the CET involves in a crucial way the use of a (arbitrary) connection providing a (arbitrary) complement of the kernel of the pre-symplectic structure.
In this work, we provide an alternative, even if equivalent, proof of the theorem. We limit ourselves to the context of vector spaces.
@@
In Section 1, we recall Oh and Park’s proof of the CET for pre-symplectic vector spaces.
In Section 2, we introduce our alternative proof of the theorem for pre-symplectic vector spaces.
@@ 1 Classical version of the Coisotropic Embedding Theorem
In this section, we recall the classical version of the CET for pre-symplectic vector spaces.
@@
Definition 1.1 (Symplectic vector space).
A symplectic vector space is a vector space $V$ equipped with a skew-symmetric, bilinear, non-degenerate form $\omega$, say $(V,\,\omega)$.
Here, we recall Darboux theorem for vector spaces which is a corollary of a more general result available for smooth differential manifolds (see [AMR88]).
Theorem 1.2 (Darboux theorem for symplectic vector spaces).
Given a symplectic vector space $(V,\,\omega)$, there exists a basis:
=$\mathcal{B}\,=\,\left\{\,\frac{\partial}{\partial x_{1}},\,...,\,\frac{\partial}{\partial x_{m}},\,\frac{\partial}{\partial y^{1}},\,...,\,\frac{\partial}{\partial y^{m}}\,\right\}\,,$
@@
$\left\{\,x^{1},\,...,\,x^{m},\,y_{1},\,...,\,y_{m}\,\right\}\,,$
such that, given its dual basis:
=$\mathcal{B}^{\star}\,=\,\left\{\,{\rm d}x_{1},\,...,\,{\rm d}x_{m},\,{\rm d}y^{1},\,...,\,{\rm d}y^{m}\,\right\}\,,$
the symplectic form $\omega$ reads:
$\omega\,=\,{\rm d}x_{j}\wedge{\rm d}y^{j}\,,$ (1)
where $\wedge$ is the skew-symmetric tensor product of elements in $\mathcal{B}^{\star}$.
Definition 1.3 (Symplectic orthogonal of a vector subspace).
Consider a symplectic vector space $(V,\,\omega)$. Consider a vector subspace $W$ of $V$. The symplectic orthogonal of $W$, say $W^{\perp}$ is defined as the set of vectors $v\in V$ such that:
[…]
Definition 1.4 (Coisotropic subspace of a symplectic vector space).
Given a symplectic vector space $(V,\,\omega)$, a vector subspace $W$ of $V$ is said to be coisotropic if $W^{\perp}\subseteq W$.
Definition 1.5 (Pre-symplectic vector space).
A pre-symplectic vector space is a vector space $V$ equipped with a skew-symmetric, bilinear form $\omega$, say $(V,\,\omega)$.
Here, we recall the analogous of Theorem 1.2 for pre-symplectic vector spaces. Again, this is a corollary of a more general result available for smooth differential manifolds [AMR88].
Theorem 1.6 (Darboux theorem for pre-symplectic vector spaces).
Given a pre-symplectic vector space $(V,\,\omega)$, there exists a basis:
=$\mathcal{B}\,=\,\left\{\,\frac{\partial}{\partial x_{1}},\,...,\,\frac{\partial}{\partial x_{m}},\frac{\partial}{\partial y^{1}}\,...,\,\frac{\partial}{\partial y^{m}},\,\frac{\partial}{\partial z^{1}},\,...,\,\frac{\partial}{\partial z^{r}}\,\right\}\,,$
where $n=2m+r$ is the dimension of $V$, $r$ being the dimension of $\mathrm{ker}\,\omega$, with the associated system of coordinates on $V$:
$\left\{\,x_{1},\,...,\,x_{m},\,y^{1},\,...,\,y^{m},\,z^{1},\,...,\,z^{r}\,\right\}\,,$
such that, given its dual basis:
=$\mathcal{B}^{\star}\,=\,\left\{\,{\rm d}x_{1},\,...,\,{\rm d}x_{m},\,{\rm d}y^{1},\,...,\,{\rm d}y^{m},\,{\rm d}z^{1},\,...,\,{\rm d}z^{r}\,\right\}\,,$
the pre-symplectic form $\omega$ reads:
$\omega\,=\,{\rm d}x_{j}\wedge{\rm d}y^{j}\,,$ (3)
and, thus, $\mathrm{ker}\,\omega$ is spanned, at each point of $V$, by:
Let us now recall the classical way of proving the coisotropic embedding theorem.
Theorem 1.7 (Coisotropic embedding theorem for vector spaces).
Let $(V,\omega)$ be a pre-symplectic vector space. There exists a symplectic vector space $(\tilde{V},\tilde{\omega})$ such that $(V,\,\omega)$ is a coisotropic subspace of $(\tilde{V},\,\tilde{\omega})$. The vector space $(\tilde{V},\,\tilde{\omega})$ is called the symplectic thickening of $(V,\,\omega)$.
@@
  • Construction of the symplectic thickening. Consider a pre-symplectic vector space $(V,\omega)$. Denote by $K$ the kernel of $\omega$, which is by assumption the same for each point of $V$. The vector space $V\times K$ is a vector subspace of $\mathbf{T}V\simeq V\times V$ (where $\simeq$ denote that the two spaces are isomorphic). Consider the cotangent bundle $\mathbf{T}^{\star}V\,\simeq V\times V^{\star}$, where $V^{\star}$ denotes the dual vector space of $V$. Consider the dual of $K$, say $K^{\star}$. Then, the symplectic thickening of $V$ is the vector space $\tilde{V}\,=\,V\times K^{\star}$ which is a vector subspace of $\mathbf{T}^{\star}V$. In the next step we prove why such a symplectic thickening is a symplectic vector space.
  • Construction of the symplectic structure. Consider a complement $W$ of $K$ into $V$, namely a vector subspace of $V$ such that $V\,=\,W\oplus K$. It is not unique. Indeed, since $K$ is spanned by:
@@
    the complement $W$ may be any of the following vector subspaces of $V$:
@@
    where ${P_{x}}^{ja}$ and ${P_{y}}_{j}^{a}$ are the entries of rectangular matrices depending on $(x,\,y,\,z)$. The vectors appearing in the latter equation, which are parametrized by the arbitrary values of ${P_{x}}^{ja}$ and ${P_{y}}_{j}^{a}$, span the kernel of the following system of linear forms on $V$:
@@
    Such a system of linear forms fulfill the conditions:
@@
    which ensures that the (1, 1)11(1,\,1)( 1 , 1 )-tensor:
@@
    is idempotent, namely it is a "projector" onto $K$. Therefore, the (1, 1)11(1,\,1)( 1 , 1 )-tensor just constructed is such that its image is $K$ and its kernel is the complementary space $W$. Such a tensor is what is usually referred to as a connection on $V$ and it is a useful way of parametrizing the arbitrariness in choosing the complement $W$.
    The decomposition $V\,=\,W\oplus K$ provided by $P$ allows us to extend $\omega$ to a symplectic form on $\tilde{V}$. Indeed, associated to $P$, there is the following linear form on $\tilde{V}$:
@@
    where $\left\{\,{p_{z}}_{a}\,\right\}_{a=1,...,r}$ is a system of coordinates on $K^{\star}$. The linear form $\theta^{P}$ evidently depends on the particular connection $P$ chosen as stressed by the notation. It allows one to define a symplectic form on $\tilde{V}$ as:
    $\tilde{\omega}\,=\,\tau^{\star}\omega+{\rm d}\theta^{P}\,,$ (10)
    where $\tau$ is the projection from $V\times K^{\star}$ to $V$. It reads:
@@
    This form is symplectic. Indeed, it is skew-symmetric and bilinear by definition and it has no kernel as it can be proven via a straightforward computation.
  • Check of the coisotropic condition. The original pre-symplectic vector space $V$ can be naturally embedded as the zero-section of $\tau$ into $\tilde{V}$, namely, it can be recovered as the subspace of $\tilde{V}$ selected by the conditions
[…]
    A direct computation shows that the symplectic orthogonal of $V$ with respect to the symplectic structure $\tilde{\omega}$ is void, namely $V$ is trivially a coisotropic subspace of $\tilde{V}$.
@@ 2 A new proof of the Coisotropic Embedding Theorem
Here, we propose our alternative proof of the theorem.
Our proof will follow the same three steps as the classical one:
  • Construction of the symplectic thickening. Given a pre-symplectic vector space $(V,\,\omega)$, the cotangent bundle $(\mathbf{T}^{\star}V,\,\omega_{\mathbf{T}^{\star}V})\,\simeq\,(V\times V^{\star},\omega_{\mathbf{T}^{\star}V})$ is a symplectic vector space, equipped with the canonical symplectic structure $\omega_{\mathbf{T}^{\star}V}$. Denote by $\rho$ the canonical projection $\rho\;\colon\;\;\mathbf{T}^{\star}V\to V$. Denote by:
@@
    ($r$ being again the dimension of $\mathrm{ker}\,\omega$ and $n=2m+r$ being the dimension of $V$) a system of Darboux coordinates on $V$ and by:
@@
    an adapted system of coordinates on $\mathbf{T}^{\star}V$. In this system of coordinates
@@
    The form $\omega^{\prime}\,=\,\omega_{\mathbf{T}^{\star}V}+\tau^{\star}\omega$ is symplectic as well. Indeed, to prove that it is non-degenerate it is sufficient to prove that
@@
    is different from zero for $n$ equal to half the dimension of $\mathbf{T}^{\star}V$. A direct computation shows that $\omega^{\prime n}\,=\,\omega_{\mathbf{T}^{\star}V}^{n}$, where $\omega_{\mathbf{T}^{\star}V}^{n}$ is different from zero because $\omega_{\mathbf{T}^{\star}V}$ is symplectic.
@@
    Recall that $\left\{\,\frac{\partial}{\partial z^{a}}\,\right\}_{a=1,...,r}$ is a system of generators of $\mathrm{ker}\,\omega\,=\,K$. Their cotangent lifts are those vector fields $\tilde{K}_{a}$ tangent to $\mathbf{T}^{\star}V$ defined by the condition
[…]
@@
    Denote by $\left\{\,H_{a}\,\right\}_{a=1,...,r}$ the Hamiltonian functions associated to $\tilde{K}_{a}$ via $\omega^{\prime}$, namely those functions staisfying
@@
    These Hamiltonian functions, together with the vectors in the original pre-symplectic vector space $V$, determine a subspace of $\mathbf{T}^{\star}V$
    $\tilde{V}\,=\,\left\{\,\frac{\partial}{\partial x_{j}},\,\frac{\partial}{\partial y^{j}},\,\frac{\partial}{\partial z^{a}},\,\frac{\partial}{\partial{p_{z}}_{a}}\,\right\}_{j=1,...,m,\,a=1,...,r}\,,$ (19)
    namely the same constructed in Section 1.
@@
    There are many, non-equivalent ways of embedding $\tilde{V}$ into $\mathbf{T}^{\star}V$. In general, an embedding of $\tilde{V}$ into $\mathbf{T}^{\star}V$ is any map of the type:
    $\begin{split}\tilde{V}\to\mathbf{T}^{\star}V\;\;:\;\;&(\,x_{j},\,y^{j},\,z^{a},\,{p_{z}}_{a}\,)\\ \mapsto&\left(\,x_{j},\,y^{j},\,z^{a},\,{p_{x}}^{j}\,=\,f^{j}(x,\,y,\,z,\,p_{z}),\,{p_{y}}_{j}\,=\,g_{j}(x,\,y,\,z,\,p_{z}),\,{p_{z}}_{a}\,\right)\,,\end{split}$ (20)
    with $f^{j}$ and $g_{j}$ smooth functions of their arguments. However, here we are going to consider only those which preserve the cotangent structure of $\mathbf{T}^{\star}V$. This class of embeddings is of the type
    $\begin{split}\mathfrak{i}\;\colon\;\;\tilde{V}\to\mathbf{T}^{\star}V\;\;:\;\;&(\,x_{j},\,y^{j},\,z^{a},\,{p_{z}}_{a}\,)\\ \mapsto&\left(\,x_{j},\,y^{j},\,z^{a},\,{p_{x}}^{j}\,=\,-{P_{x}}^{ja}{p_{z}}_{a},\,{p_{y}}_{j}\,=\,-{P_{y}}^{a}_{j}{p_{z}}_{a},\,{p_{z}}_{a}\,\right)\,,\end{split}$ (21)
@@
    where ${P_{x}}^{ja}$ and ${P_{y}}^{a}_{j}$ are the entries of two rectangular matrices of maximal rank depending on $(x,\,y,\,z)$ and where the minus sign is an arbitrary choice. The fact that $\mathfrak{i}$ is linear in the ${p_{z}}_{a}$ ensures that $\mathfrak{i}$ preserves the zero-section of $\rho$, namely $\rho\bigr{|}_{Im\,\mathfrak{i}}$ is still a submersion onto $V$. On the other hand, the maximal rank hypothesis ensures that the fibres of $\rho\bigr{|}_{Im\,\mathfrak{i}}$ are all diffeomorphic, and, thus, the image of $\mathfrak{i}$ defines a vector subbundle of $\mathbf{T}^{\star}V$.

### v3

# The coisotropic embedding theorem for pre-symplectic manifolds: an alternative proof

@@
Abstract
We present an alternative proof of the Coisotropic Embedding Theorem in which the geometric choice of a connection is recast as the algebraic choice of an embedding into the cotangent bundle. The symplectic thickening is then identified as the submanifold determined by the Hamiltonian momenta conjugate to the kernel directions of the pre-symplectic form.
@@ Acknowledgements
I acknowledge financial support from Next Generation EU through the project 2022XZSAFN – PRIN2022 CUP: E53D23005970006. I am a member of GNSAGA (Indam).
@@ Introduction
The Coisotropic Embedding Theorem (CET), originally proven by Mark Gotay in his seminal work on pre-symplectic manifolds Got (82), establishes that any pre-symplectic manifold can be embedded into a symplectic manifold in such a way that the original pre-symplectic manifold becomes a coisotropic submanifold. It has significant applications in both theoretical mathematics and applied fields, where it facilitates the study of constrained mechanical systems and reduction theory from various perspectives, as we will recall in this introduction.
@@
In mathematical physics, Hamiltonian systems are geometrically formulated by means of symplectic manifolds (usually the cotangent bundle of a differential manifold, sometimes referred to as Phase Space) endowed with a Hamiltonian function Arn (89); AMRC (78). On the other hand, constrained Hamiltonian systems are usually described on submanifolds of the Phase Space (selected by the constraints) on which the symplectic form becomes degenerate (see GNH (78); CDI+ (24); CDI+22a and references therein). Among various interesting physical systems, all gauge theories lie in this class of dynamical systems. The CET ensures that these constrained systems can still be analyzed within the framework of symplectic geometry by embedding the presymplectic manifold into a larger symplectic manifold Got (82); AMRC (78); GS (90).
@@
Recent advances in mathematical physics have highlighted the utility of the CET in various contexts. For instance, in CDI+22a , the theorem is used to establish a Noether-like one-to-one correspondence between symmetries and constants of motion within pre-symplectic Hamiltonian systems. This correspondence is fundamental in understanding the conservation laws and invariant properties of constrained mechanical systems, thereby extending Noether’s classical theorem to a broader context. Furthermore, the CET has been employed to provide a Poisson structure on the space of solutions of certain field theories. For example, in CDI+ (24); CDI+22b ; CDI+ (23), the theorem is used to pass from a pre-symplectic framework to a symplectic one, enabling the application of symplectic geometry and Poisson structures to the field theories considered, gauge theories, and Palatini’s formulation of General Relativity in particular. Moreover, the theorem has been instrumental in addressing the inverse problem of the calculus of variations. In recent works Sch24b ; Sch24a , the CET has been used to provide a solution for the inverse problem for a class of implicit differential equations. These examples illustrate how the theorem can be leveraged to construct variational principles for systems that are not explicitly variational. Additionally, it is worth mentioning that the CET has significant applications within the Geometric Quantization program GS (81). Geometric Quantization is a procedure for constructing a quantum theory from a classical mechanical system, and requires a symplectic structure for the classical system as starting point. By embedding a pre-symplectic manifold into a symplectic one, the theorem facilitates the Quantization of constrained systems by ensuring that the necessary symplectic structure is available.
@@
Several reformulations of the theorem have been proposed over the years. We mention the approach of V. Guillemin and S. Sternberg that proved an equivariant (with respect to the action of a Lie group) version of the theorem GS (90) and the approach of Y. G. Oh and J. S. Park OP (05) applied to the deformation of coisotropic submanifolds.
In this work, we provide an alternative, albeit equivalent, proof of the theorem. While the classical proof of the CET is well-established, this work aims to present an alternative construction that offers a distinct conceptual and pedagogical viewpoint. Instead of geometrically constructing a complement to the kernel of the pre-symplectic form, our approach is rooted in the canonical structure of the cotangent bundle $\mathbf{T}^{\star}M$. The symplectic thickening is identified by selecting the submanifold of $\mathbf{T}^{\star}M$ determined by the Hamiltonian functions conjugate to the kernel directions of the pre-symplectic structure. This perspective recasts the usual arbitrary choice of a connection—a purely geometric object—into the choice of an embedding into the phase space, a concept more directly tied to its algebraic structure.
@@
In Section˜1, we recall Oh and Park’s proof of the CET.
In Section˜2, we introduce our alternative proof of the theorem.
@@ 1 Classical version of the Coisotropic Embedding Theorem
In this section, we recall the classical version of the CET.
@@
Definition 1.1 (Symplectic manifold).
A symplectic manifold is a manifold $M$ equipped with a closed, non-degenerate $2$-form $\omega$, say $(M,\,\omega)$.
Here, we recall Darboux theorem for symplectic manifolds (see, for instance, AMR (88)).
Theorem 1.2 (Darboux theorem for symplectic manifolds).
Given a $2m$-dimensional symplectic manifold $(M,\,\omega)$, there exists a system of local coordinates
$\left\{\,x_{1},\,...,\,x_{m},\,y^{1},\,...,\,y^{m}\,\right\}\,$,
such that the symplectic form $\omega$ locally reads:
$\omega\,=\,{\rm d}x_{j}\wedge{\rm d}y^{j}\,.$ (1)
Definition 1.3 (Symplectic orthogonal of a submanifold).
Consider a symplectic manifold $(M,\,\omega)$. Consider a submanifold $N$ of $M$. The symplectic orthogonal of $T_{n}N$, say $T_{n}N$$^{\perp}$ is defined as the set of vectors $v\in T_{n}M$ such that:
$\omega_{n}(v,\,w)\,=\,0\,,\;\;\;\forall\,\,w\in T_{n}N\,.$ (2)
Definition 1.4 (Coisotropic submanifold of a symplectic manifold).
Given a symplectic manifold $(M,\,\omega)$, a submanifold $N$ of $M$ is said to be coisotropic if ${T_{n}N}^{\perp}\subseteq T_{n}M\,,\;\;\;\forall\,\,n\in N$.
Definition 1.5 (Pre-symplectic manifold).
A pre-symplectic manifold is a $n$-dimensional manifold $M$ equipped with closed, non-necessarily non-degenerate, $2$-form $\omega$, say $(M,\,\omega)$.
Here, we recall the analog of Theorem˜1.2 for pre-symplectic manifolds (see, for instance, AMR (88)).
Theorem 1.6 (Darboux theorem for pre-symplectic manifolds).
Given a $(2m+r)$-dimensional pre-symplectic manifold $(M,\,\omega)$, there exists a system of local coordinates on $M$
$\left\{\,x_{1},\,...,\,x_{m},\,y^{1},\,...,\,y^{m},\,z^{1},\,...,\,z^{r}\,\right\}\,$,
such that the pre-symplectic form $\omega$ locally reads:
$\omega\,=\,{\rm d}x_{j}\wedge{\rm d}y^{j}\,$, (3)
and, thus, $\mathrm{ker}\,\omega$ is spanned, at each point, by:
Let us now recall the classical way of proving the CET.
Theorem 1.7 (Coisotropic embedding theorem).
@@
  • Construction of the symplectic thickening. Consider a pre-symplectic manifold $(M,\omega)$. Denote by $K$ the kernel of $\omega$, which is, by assumption, the same for each point of $M$. Thus, it defines a vector subbundle $\mathbf{K}$ of $\operatorname{\textup{{T}}}M$, called the characteristic bundle. Consider the dual vector bundle $\mathbf{K}^{\star}$, $\tau$ denoting the canonical projection onto $M$. The symplectic thickening of $M$ is (a suitable submanifold of) such a vector bundle.
  • Construction of the symplectic structure. Consider a complement $W$ of $K$ into $T_{m}M$, namely a vector subspace of $T_{m}M$ such that $T_{m}M\,=\,W\oplus K$. It is not unique. Indeed, since $K$ is spanned by:
@@
    the complement $W$ may be any of the following vector subspaces of $T_{m}M$:
@@
    where ${P_{x}}^{ja}$ and ${P_{y}}_{j}^{a}$ are the entries of real-valued rectangular matrices of maximal rank depending on the point $m$. The vectors appearing in the latter equation, which are parametrized by the arbitrary values of ${P_{x}}^{ja}$ and ${P_{y}}_{j}^{a}$, span the kernel of the following system of $1$-forms on $M$:
@@
    Such a system of $1$-forms fulfills the conditions:
@@
    which ensures that the (1, 1)(1,\,1)-tensor field:
@@
    is idempotent, namely, it is, at each point, a "projector" onto $K$. Therefore, the (1, 1)(1,\,1)-tensor just constructed is such that, at each point, its image is $K$ and its kernel is the complementary space $W$. Such a tensor field is what is usually referred to as a connection on $M$, and it is a useful way of parametrizing the arbitrariness in choosing the complement $W$.
    The decomposition $T_{m}M\,=\,W\oplus K$ provided by $P$ allows us to extend $\omega$ to a symplectic form on $\tilde{M}$. Indeed, associated to $P$, there is the following $1$-form on $\mathbf{K}^{\star}$:
    $\theta^{P}_{p}(X_{p})\,=\,\langle\,p,\,P\circ T\tau(X_{p})\,\rangle\,$, (9)
    where $X_{p}\in\mathbf{T}_{p}\mathbf{K}^{\star}$, and $p$ has to be interpreted as a point in $\mathbf{K}^{\star}$ on the left-hand side, and as a covector on $K_{m}$, with $m=\tau(p)$, on the right-hand side. Locally, the latter $1$-form reads
@@
    where $\left\{\,{p_{z}}_{a}\,\right\}_{a=1,...,r}$ is a system of coordinates on the fibers of $\tau$. The $1$-form $\theta^{P}$ evidently depends on the particular connection $P$ chosen, as stressed by the notation. It allows one to define the following $2$-form on $\mathbf{K}^{\star}$
    $\tilde{\omega}\,=\,\tau^{\star}\omega+{\rm d}\theta^{P}\,.$ (11)
    It locally reads:
@@
    The $2$-form (12) is closed by construction, and, as it can be proven via a straightforward computation, is non-degenerate in a tubular neighborhood of the zero-section of $\tau$, which will be denoted as $\tilde{M}$ and represents the symplectic manifold we are searching for. As it is shown in CDI+22b , $\tilde{M}$ coincides with the whole $\mathbf{K}^{\star}$ when the connection $P$ is flat.
  • Check of the coisotropic condition. The original pre-symplectic manifold $M$ can be naturally embedded as the zero-section of $\tau$ into $\tilde{M}$, namely, it can be recovered as the submanifold of $\tilde{M}$ locally selected by the conditions
    ${p_{z}}_{a}\,=\,0\,,\;\;\;\forall\,\,a=1,...,r\,.$
    A direct computation shows that the symplectic orthogonal of $M$ with respect to the symplectic structure $\tilde{\omega}$ is $\operatorname{\textup{{T}}}_{m}M^{\perp}=K$. Thus, $\operatorname{\textup{{T}}}_{m}M^{\perp}\,=\,K\subset T_{m}M$ is a coisotropic subspace of $\tilde{M}$ by definition.
@@ 2 A new proof of the Coisotropic Embedding Theorem
Here, we propose our alternative proof of the theorem for pre-symplectic manifolds. Our proof will follow the same three steps as the classical one.
  • Construction of the symplectic thickening. Given the pre-symplectic manifold $(M,\omega)$, we begin with its cotangent bundle, $(\operatorname{\textup{{T}}}^{\star}M,\,\omega_{\operatorname{\textup{{T}}}^{\star}M})$, which is a canonical symplectic manifold. Denote by $\rho$ the canonical projection $\rho\;\colon\;\;\mathbf{T}^{\star}M\to M$. Denote by:
@@
    ($r$ being again the dimension of $\mathrm{ker}\,\omega$ and $n=2m+r$ being the dimension of $M$) a system of Darboux coordinates on $M$ and by:
@@
    an adapted system of coordinates on $\mathbf{T}^{\star}M$. In this system of coordinates
@@
    The form $\omega^{\prime}\,=\,\omega_{\mathbf{T}^{\star}M}+\rho^{\star}\omega$ is symplectic as well. Indeed, to prove that it is non-degenerate it is sufficient to prove that
@@
    is different from zero for $n$ equal to half the dimension of $\mathbf{T}^{\star}M$. A direct computation shows that $\omega^{\prime n}\,=\,\omega_{\mathbf{T}^{\star}M}^{n}$, where $\omega_{\mathbf{T}^{\star}M}^{n}$ is different from zero because $\omega_{\mathbf{T}^{\star}M}$ is symplectic.
@@
    Recall that $\left\{\,\frac{\partial}{\partial z^{a}}\,\right\}_{a=1,...,r}$ is a system of generators of $\mathrm{ker}\,\omega\,=\,K$. Their cotangent lifts are those vector fields $\tilde{K}_{a}$ tangent to $\mathbf{T}^{\star}M$ defined by the condition
    $\mathcal{L}_{\tilde{K}_{a}}\omega_{\mathbf{T}^{\star}M}\,=\,0\,$, (16)
@@
    Denote by $\left\{\,H_{a}\,\right\}_{a=1,...,r}$ the Hamiltonian functions associated to $\tilde{K}_{a}$ via $\omega^{\prime}$, namely those functions satisfying
@@
    The Hamiltonian functions $H_{a}$ are linear on the fibers of the cotangent bundle and represent the momenta conjugate to the kernel directions $z_{a}$.
    We now construct the thickening manifold $\tilde{M}$ by applying the Gel’fand-Kolmogorov theorem MM (94).
    Proposition 2.1.
    The bundle $\mathbf{K}^{\star}$ of the classical proof of the CET is the manifold associated, by means of the Gel’fand-Kolmogorov theorem, with the algebra $\mathcal{A}\subset C^{\infty}(\operatorname{\textup{{T}}}^{\star}M)$ generated by functions pulled back from $M$ and the kernel-momenta functions $\{H_{a}\}$.
    Proof.
    A point in the manifold associated with the algebra $\mathcal{A}$ can be identified with a character $\chi\,:\;\mathcal{A}\to\mathbb{R}$, which is an algebra homomorphism. Any such character is determined by its action on the generators of $\mathcal{A}$.
    1. 1.
      The restriction of $\chi$ to the subalgebra of functions pulled back from the base, $\{f\circ\rho\;:\;\;f\in C^{\infty}(M)\}$, is a character on $C^{\infty}(M)$. The space of characters of $C^{\infty}(M)$ is the manifold $M$ itself. Therefore, for each $\chi$, there exists a unique point $m\in M$ such that $\chi(f\circ\rho)=f(m)$ for all $f\in C^{\infty}(M)$. This fixes the base point of our new manifold.
    2. 2.
      The action of $\chi$ on each Hamiltonian generator $H_{a}$ yields a real number, $p_{a}:=\chi(H_{a})$. This set of numbers $\{p_{a}\}_{a=1}^{r}$ can be used to define a covector on the kernel fiber $K_{m}$. Specifically, we define $\alpha_{m}\in K_{m}^{\star}$ as the linear functional on $K_{m}$ such that $\braket{\alpha_{m},\,\frac{\partial}{\partial z^{a}}}=p_{a}$.
    Thus, each character $\chi$ corresponds to a unique pair $(m,\alpha_{m})$ where $m\in M$ and $\alpha_{m}\in K_{m}^{\star}$. This establishes a bijection between the space of characters of $\mathcal{A}$ and the total space of the dual bundle $K^{\star}$. Since $K^{\star}$ is a vector bundle over a smooth manifold, it is itself a smooth manifold. ∎
@@
    There are many, non-equivalent ways of embedding $\mathbf{K}^{\star}$ into $\mathbf{T}^{\star}M$ as a vector subbundle. This class of embeddings is of the type
[…]
@@
    where ${P_{x}}^{ja}$ and ${P_{y}}^{a}_{j}$ are the entries of two rectangular matrices of maximal rank depending on the base coordinates $(x_{j},\,y^{j},\,z^{a})$, and where the minus sign is an arbitrary choice. The fact that $\mathfrak{i}$ is linear in the ${p_{z}}_{a}$ ensures that $\mathfrak{i}$ preserves the zero-section of $\rho$. On the other hand, the maximal rank hypothesis ensures that the fibres of $\rho\bigr{|}_{Im\,\mathfrak{i}}$ are all diffeomorphic, and, thus, the image of $\mathfrak{i}$ defines a vector subbundle of $\mathbf{T}^{\star}M$.
    It is crucial to note that the functions ${P_{x}}^{ja}$ and ${P_{y}}^{a}_{j}$ that define the embedding $\mathfrak{i}$ play precisely the same role as the components of the connection $1$-forms (6) in the classical proof. Our "choice of embedding" is therefore the direct counterpart of the classical "choice of a connection".